OpenStudy (anonymous):

Solve expression: 9/5*6 + 32?

OpenStudy (anonymous):

@AccessDenied @mathsss @ilikephysics2

OpenStudy (anonymous):

It's a bit ambiguous \[\frac{9}{5 \cdot 6} + 32 \] Or \[\frac{9}{5} \cdot 6 + 32 \]

OpenStudy (anonymous):

Second one.

OpenStudy (anonymous):

Okay http://www.wolframalpha.com/input/?i=%289%2F5%29*6+%2B+32 This tool will be more helpful to you. Converting from Celsius to Fahrenheit, are we?

OpenStudy (anonymous):

42.8?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

42.8 is correct